OpenStudy (anonymous):

3^-2+2^2/4^-2/2^-2 evaluate

OpenStudy (anonymous):

\[3^{-2}+2^{-2}\div 4^{-2} + 2^{-2}\]

OpenStudy (anonymous):

sorry 3^-2 + 2^2

OpenStudy (anonymous):

\[\large \frac{3^{-2}+2^2}{4^{-2}+2^{-2}}\]?

OpenStudy (anonymous):

\[\frac{ 3^{-2}+2^{-2} }{ 4^{-2}+2^{-2} }\]

OpenStudy (anonymous):

yeah! :) @satellite73

OpenStudy (anonymous):

there is not short cut here you have to compute each number separately

OpenStudy (anonymous):

for example \[3^{-2}=\frac{1}{3^2}=\frac{1}{9}\]

OpenStudy (anonymous):

\[\huge \frac{ \frac{ 1 }{ 3^2 }+\frac{ 1 }{ 2^2 } }{ \frac{ 1 }{ 4^2 }+\frac{ 1 }{ 2^2 } }\]

OpenStudy (anonymous):

\[2^2=4\\ 4^{-2}=\frac{1}{4^2}=\frac{1}{16}\\ 2^{-2}=\frac{1}{4}\]

OpenStudy (anonymous):

Should be pretty simple from there :P

OpenStudy (anonymous):



OpenStudy (anonymous):

@iambatman i think i might be \[\huge \frac{ \frac{ 1 }{ 3^2 }+2^2 }{ \frac{ 1 }{ 4^2 }+\frac{ 1 }{ 2^2 } }\]

OpenStudy (anonymous):

do i just divide them now?

OpenStudy (anonymous):

Ohhhh yeah nice catch satellite

OpenStudy (anonymous):

you have to do the annoying arithmetic top and bottom separately, then divide by multiplying by the reciprocal of the denominator

OpenStudy (anonymous):

sorry what do you mean by top and bottom seperatly?

OpenStudy (anonymous):

\[\huge \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } = \frac{ ad }{ bc }\]

OpenStudy (anonymous):

for example \[\frac{1}{9}+4=\frac{37}{9}\] that is the top

OpenStudy (anonymous):

i got 592/45 as my final answer is that right ? @satellite73 @iambatman

OpenStudy (anonymous):